Solve and show steps to answer
15/x- 15/x-2 = -2

Answer by guapa(62)

15/x - 15/(x-2)= -2 First we have to make a restriction to avoid division by 0.
x cannot be 0 or 2.
We nee to find the LCD to clear the equation of all fractions. In this case the LCD is x(x-2)
So we get: 15(x-2)- 15x= -2x(x-2)
15x-30-15x=-2x²+4x
2x²-4x-30=0
x²-2x-15=0
x²-5x+3x-15=0
x(x-5)+3(x-5)=0
(x+3)(x-5)=0
x+3=0, x=-3
x-5=0, x=5
The solution set is {-3,5}
Both solutions are possible since we made the restriction of 0 and 2.